On March 1, 2014, Newton Company purchased land for an office site by paying $1,800,000 cash. Newton began construction on the office building on March 1. The following expenditures were incurred for construction:

Date                                                   Expenditures

March 1, 2014                                  $   1,200,000

April 1, 2014                                          1,680,000

May 1, 2014                                          3,000,000

June 1, 2014                                         4,800,000

The office was completed and ready for occupancy on July 1. To help pay for construction, and purchase of land $2,400,000 was borrowed on March 1, 2014 on a 9%, 3-year note payable. Other than the construction note, the only debt outstanding during 2014 was a $1,000,000, 12%, 6-year note payable dated January 1, 2014.

a) The weighted-average accumulated expenditures on the construction project during 2014 were

b) The actual interest cost incurred during 2014 was

c) Assume the weighted-average accumulated expenditures for the construction project are $2,900,000. The amount of interest cost to be capitalized during 2014 is

Please provide work
Answer #1

a) Weighted-average accumulated expenditures on the construction project during 2014=

($3,000,000 x 4/12) + (1,680,000 x 3/12) + (3,000,000 x 2/12) + (4,800,000 x 1/12) = $2,320,000

b) Actual interest cost incurred during 2014

($2,400,000 x 9% x 10/12) + ($1,000,000 x 12%) = $300,000

c)  amount of interest cost to be capitalized during 2014

=($2,400,000 × 9%) + ($500,000 ×12%) = $276,000.

Image of page 4